# Realistic Environmental Injection Attacks on GUI Agents

### v1

@@ 6.3. RQ2: Generalization Ability Across LVLMs
In this RQ, we evaluate the generalization ability of our trigger image across multiple LVLMs. Specifically, we aim to examine whether a trigger trained on one model can generalize to other unseen models in a black-box setting.
@@ 6.4. RQ3: Ablation Study
Table 3. ASRs (%) under different ablation settings. NetEase is used as an abbreviation for NetEase Cloud Music. The percentage shown at the top-right of each website name denotes the trigger coverage ratio (percentage of screenshot pixels occupied by the trigger image).
@@
Amazon4.67% Taobao7.59% RedNote8.19% Bilibili9.97% NetEase8.17% QQ Music5.99%
@@
Analyses.Contribution of LES. Removing LES leads to a consistent reduction in ASR across all websites. For example, with OS-Atlas-Base-7B on NetEase Cloud Music, the ASR decreases from 43.75% to 20.11% when LES is removed. Similarly, with LLaVA-1.5-13B on RedNote, the ASR drops from 60.75% to 34.42%. These results indicate that large-scale automatic simulation of realistic and diverse visual context during training is critical to achieving effective attacks. ❷ Contribution of ABH. Removing ABH also reduces ASR on every website. Notably, the largest declines appear when the trigger image occupies a smaller portion of the screenshot. For instance, with OS-Atlas-Base-7B on QQ Music, where the trigger covers 5.99% of the screenshot, the ASR decreases by 6.32%, while on RedNote, with a coverage of 8.19%, the decrease is only 2.32%. For LLaVA-1.5-13B, the ASR drops by 10.34% on Amazon, where the trigger coverage is 4.67%, but only 4.12% on Bilibili, where the coverage reaches 9.97%. We hypothesize that this is because, when the trigger image occupies a larger portion of the screenshot, the model can more readily focus on the trigger image region even without explicit guidance.
@@
7. Case Study
Figure 5. Case study: GUI agent’s behavior for the task “open the first playlist in the homepage recommendations of NetEase Cloud Music.” (a) Homepage with the uploaded trigger image. (b) Without attack: correct navigation to the first playlist. (c)With attack: incorrect navigation to a promotional site.
@@ 8. Discussion
8.1. Defenses
We investigate several practical defenses for GUI agents under the threat model considered in this paper and evaluate their ability to reduce the effectiveness of Chameleon. Experiments are conducted on LLaVA-1.5-13B and OS-Atlas-Base-7B using the NetEase Cloud Music and RedNote as target websites. Effectiveness is measured by the Attack Success Rate (ASR), and the results are shown in Figure 6.
Safety prompt. Safety prompts (Wang et al., 2024c; Zhang et al., 2025b) are a widely adopted technique for guiding models away from unsafe behavior. We insert a safety prompt into the system prompt that explicitly directs the agent to follow the user’s command strictly and to disregard any potentially malicious content in the observation. The specific prompt used is shown below:
However, the ASR shows almost no reduction. For example, in the case of a LLaVA-1.5-13B powered agent on RedNote, the ASR decreases only marginally from 60.75% to 60.21%. This indicates that safety prompts alone are insufficient to defend against Chameleon.
Verifier. Another common defense is to verify model outputs before execution. We employ DeepSeek-V3 (Liu et al., 2024) as an external verifier that examines each candidate action together with the user instruction and aborts execution if the action appears inconsistent or risky. For the OS-Atlas-Base-7B powered agent on NetEase Cloud Music, this verification reduces the ASR from 43.75% to 36.42%. However, we also observe false positives where benign actions are blocked, which reduces overall utility. Furthermore, this approach introduces additional latency and inference cost since it relies on a secondary model.
Random noise on uploaded images. Since Chameleon performs attacks by applying adversarial perturbations, one potential countermeasure is to add random noise to uploaded images. We assume that a website administrator automatically adds noise $\eta$ with an $\ell_{\infty}$ bound $|\eta|_{\infty}\leq\epsilon$. With $\epsilon=\tfrac{8}{255}$, this defense proves highly effective: on RedNote, the ASR of a LLaVA-1.5-13B powered agent drops to 11.84%, and for OS-Atlas-Base-7B the ASR is reduced to nearly zero. Although effective, even small bounded noise noticeably degrades image quality, which may harm user experience, particularly in scenarios where high visual fidelity is required, such as photography or digital art. As shown in Figure 7, applying random noise leads to clear quality degradation.
@@ 8.2. Threats to Validity
Trigger Images. For each website, we randomly select a single image from the crawled corpus to serve as the trigger. The choice of trigger likely affects attack effectiveness due to factors such as visual salience, color palette, and composition. Our current results therefore reflect the performance of one draw per site. Future work will expand this evaluation by considering multiple triggers for each site, analyzing sensitivity to different visual attributes, and developing principled strategies for selecting images that maximize attack effectiveness. A more systematic study of trigger characteristics would also provide insights into which types of content are more likely to compromise GUI agents in practice.
Datasets. Existing GUI agent datasets often present webpages where text and images are static or where images occupy a disproportionately large fraction of the webpage, which misaligns with real-world internet scenarios. To address this, we construct test sets using our LLM-Driven Environment Simulation, which enables large-scale generation of realistic and diverse visual contexts. Moreover, to eliminate potential data leakage, we ensure that the screenshots used in the training, validation, and test sets do not share any uploaded images. This separation guarantees that performance improvements cannot be attributed to memorization of specific samples.
@@
Ethics Statement. In this work, we propose a more practical and realistic threat model and an effective attack approach. However, our goal is not to promote malicious behavior, but rather to reveal the vulnerabilities of widely used GUI agents when deployed in real-world, open-ended internet environments. By exposing these vulnerabilities, we aim to raise awareness of the potential risks and emphasize the urgent need for robust and practical defenses. We hope that our findings will inform future research on building safer and more trustworthy web-based agent systems.
@@ 9. Conclusion
In this paper, we present a more realistic threat model in which the attacker is a regular user who can only upload a small trigger image that appears within a dynamically changing environment. To address the challenges posed by dynamic visual contexts and limited screenshot coverage, we propose Chameleon , a novel attack framework that introduces two key novelties:LLM-Driven Environment Simulation, which enables large-scale automatic generation of realistic and diverse webpage simulations, and Attention Black Hole, which explicitly guides the agent’s focus toward the trigger region. Extensive experiments across multiple websites and models demonstrate that Chameleon significantly outperforms existing methods in attack success rate. Ablation and closed-loop evaluations further confirm the effectiveness and real-world applicability of the proposed techniques. Through an evaluation of several commonly used defense strategies, we find that existing defense strategies fail to effectively mitigate the threat of Chameleon without compromising the utility of the agent. Overall, our study reveals the inherent vulnerabilities in widely used LVLM-powered GUI agents. Future work may explore automated trigger detection and effective defenses that preserve usability in open-world web environments.
@@ 10. Data Availability
The source code and access details for the datasets can be accessed through an anonymous repository at: https://github.com/zhangyitonggg/attack2gui.

### v2

@@ 5.3. RQ2: Generalization Ability Across LVLMs
In this RQ, we evaluate the generalization ability of our trigger image across multiple LVLMs. Specifically, we aim to examine whether a trigger trained on one LVLM-powered GUI agent can generalize to other unseen agents in a black-box setting.
@@ 5.4. RQ3: Ablation Study
Table 3. ASRs (%) under different ablation settings. NetEase is used as an abbreviation for NetEase Cloud Music.
@@
Amazon Taobao RedNote Bilibili NetEase QQ Music
@@
Analyses.Contribution of LES. Removing LES leads to a consistent reduction in ASR across all websites. For example, with OS-Atlas-Base-7B on NetEase Cloud Music, the ASR decreases from 43.75% to 20.11% when LES is removed. Similarly, with LLaVA-1.5-13B on RedNote, the ASR drops from 60.75% to 34.42%. These results indicate that large-scale automatic simulation of realistic and diverse visual contexts during training is critical for achieving effective attacks. ❷ Contribution of ABH. Removing ABH also reduces ASR on every website, demonstrating that explicit attention supervision is consistently beneficial. For instance, with OS-Atlas-Base-7B on QQ Music, the ASR decreases by 6.32%. For LLaVA-1.5-13B, removing ABH reduces ASR by 10.34% on Amazon and by 4.12% on Bilibili. We hypothesize that ABH improves robustness by suppressing interference from non-trigger regions and encouraging the model to maintain stable focus on the trigger across dynamically changing environments.
@@
5.5. RQ4: Possible Defenses
In this RQ, we explore practical defense strategies for GUI agents under the dynamic-environment threat model and evaluate their effectiveness in reducing the attack success rate of Chameleon.
Setup. We evaluate defenses on two GUI agents, LLaVA-1.5-13B and OS-Atlas-Base-7B, using NetEase Cloud Music and RedNote as target websites. Defense effectiveness is measured by ASR. Figure 6 summarizes the results.
5.5.1. Safety Prompt
Safety prompts (Wang et al., 2024c; Zhang et al., 2025b) are widely used to steer models away from unsafe behavior. We prepend the following instruction to the system prompt to encourage the agent to follow the user request and ignore potentially malicious content in the observation:
This defense yields almost no reduction in ASR. For example, on RedNote with a LLaVA-1.5-13B powered agent, ASR decreases only from 60.75% to 60.21%. This suggests that prompt-based steering alone is insufficient against Chameleon.
5.5.2. Verifier
A common mitigation is to verify model outputs before execution. We use DeepSeek-V3 (Liu et al., 2024) as an external verifier that checks each candidate action against the user instruction and aborts execution if the action appears inconsistent or risky. Verification reduces ASR in some settings, for example from 43.75% to 36.42% for OS-Atlas-Base-7B on NetEase Cloud Music. However, we also observe false positives that block benign actions, reducing overall utility. Moreover, this approach introduces additional latency and inference cost because it requires a secondary model call at each step.
5.5.3. Random Noise on Uploaded Images
Since Chameleon performs attacks by applying adversarial perturbations, one potential countermeasure is to add random noise to uploaded images. We assume that a website administrator automatically adds noise $\eta$ with an $\ell_{\infty}$ bound $|\eta|_{\infty}\leq\epsilon$. With $\epsilon=\tfrac{8}{255}$, this defense proves highly effective: on RedNote, the ASR of a LLaVA-1.5-13B powered agent drops to 11.84%, and for OS-Atlas-Base-7B the ASR is reduced to nearly zero. Although effective, even small bounded noise noticeably degrades image quality, which may harm user experience, particularly in scenarios where high visual fidelity is required, such as photography or digital art. As shown in Figure 7, applying random noise leads to clear quality degradation.
Answer to RQ4: Many existing defenses struggle to mitigate Chameleon without substantially degrading user experience. Safety prompts provide negligible protection, output verification offers only moderate gains with added latency and utility loss, and randomized image noise can be effective but noticeably harms visual quality.
@@ 6. Discussion
6.1. Case Study
Figure 8. Case study: GUI agent’s behavior for the task “open the first playlist in the homepage recommendations of NetEase Cloud Music.”
@@ 6.2. Threats to Validity
We identify four primary threats to the validity of our study:
@@
Datasets. Existing GUI agent datasets often present webpages where text and images are static, which misaligns with real-world internet scenarios. To address this, we construct test sets using our LLM-Driven Environment Simulation, which enables large-scale generation of realistic and diverse visual contexts. Moreover, to eliminate potential data leakage, we ensure that the screenshots used in the training, validation, and test sets do not share any uploaded images. This separation guarantees that performance improvements cannot be attributed to memorization of specific samples.
@@
Ethics Statement. In this work, we propose a more practical and realistic threat model and an effective attack approach. However, our goal is not to promote malicious behavior, but rather to reveal the hidden vulnerabilities of widely used GUI agents when deployed in real-world, open-ended internet environments. By exposing these vulnerabilities, we aim to raise awareness of the potential risks and emphasize the urgent need for robust and practical defenses. We hope that our findings will inform future research on building safer and more trustworthy web-based agent systems.
@@ 7. Conclusion
In this paper, we present a more realistic threat model in which the attacker is a regular user who can only upload trigger images that appear within a dynamically changing environment. To address the challenges posed by dynamic environment, we propose Chameleon, a novel attack framework that introduces two key novelties:LLM-Driven Environment Simulation, which enables large-scale automatic generation of realistic and diverse webpage simulations, and Attention Black Hole, which explicitly guides the agent’s focus toward the trigger region. Extensive experiments across multiple websites and models demonstrate that Chameleon significantly outperforms existing methods in attack success rate. Ablation and closed-loop evaluations further confirm the effectiveness and real-world applicability of the proposed techniques. Through an evaluation of several commonly used defense strategies, we find that existing defense strategies fail to effectively mitigate the threat of Chameleon without compromising the utility of the agent. Overall, our study reveals the inherent vulnerabilities in widely used LVLM-powered GUI agents. Future work may explore automated trigger detection and effective defenses that preserve usability in open-world web environments.
@@ 8. Data Availability
The source code and access details for the datasets can be accessed at: https://github.com/zhangyitonggg/attack2gui.